OpenStudy (anonymous):

1. A ball rolls with a constant acceleration of a = 2 m/s2 over 5 s and an initial velocity of 0 m/s. Calculate the ball’s distance from the starting point at 1.0 s intervals. Make a position-time graph for the object’s motion. In your response, show what you are given, the equation that you used, any algebra required, a table of data, and your graph.

OpenStudy (anonymous):

Since the ball is accelerating, use the following equation to determine the distance it moves. d = vi * t + ½ * a * t^2, vi = 0, a = 2 d = ½ * 2 * t^2 = t^2 d0 = 0 d1 = 1^2 = 1 d2 = 2^2 = 4 d3 = 3^2 = 9 d4 = 4^2 = 16 d5 = 5^2 = 25 When you graph the data, the graph will look like one half of a parabola. If you go to the website below you will see a graph like this one. http://www.physicsclassroom.com/class/1D...

OpenStudy (anonymous):

it wont let me go to that page..

OpenStudy (anonymous):

@pitamar

OpenStudy (anonymous):

@tyrone24 Go on google and search physics classroom/classroom/1D